LINE PARAMETERS OF CO_{2} BETWEEN 4500 AND 7000 cm^{−1}.

\underline{D. CHRIS BENNER}, V. MALATHY DEVI, Department of Physics, The College of William and Mary, Williamsburg, VA 23187-8795; LINDA R. BROWN, CHARLES E. MILLER and ROBERT A. TOTH, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, California 91109.

To enable column CO_{2} measurements from space with 0.3% precisions by the Orbital Carbon Observatory, positions and intensities were measured for 125 near-IR CO_{2} bands using Voigt line shapes. Self- and air-broadening coefficients were obtained for the stronger bands as well. In addition, measurement of the 30012 - 00001 band at 6348 cm^{−1} was repeated using constrained multispectrum fitting^{a} employing a speed dependent Voigt profile and the off-diagonal relaxation matrix elements (RME) formulation for line mixing. For this, high resolution, high signal-to-noise ratio spectra were recorded at room temperature with the McMath-Pierce Fourier transform spectrometer on Kitt Peak.

aD. Chris Benner et al. JQSRT, \underline{53}, 705-721, 1995. Benner et al. in preparation.